# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 4.2 Compiled self-test for tilted-CHSH inequalities
We now proceed to study the relations imposed by the second SOS polynomial, which imply $\hat{E}_{1}\leq\delta_{0}/\lambda_{\theta,\phi}^{2}$. Before we begin, we note that in any earlier claim that did not include a factor of $(-1)^{\mathsf{Dec}(\alpha)}$, we can always apply Lemma 3.1 to exchange the expectation over encryptions of $x=0$ to an expectation over encryptions of $x=1$ at the cost of adding $\delta_{\mathsf{QHE}}$ to the bound. This follows from the fact that such claims are bounding the expectation of operators $\mathcal{N}$ constructed from sums of products of $B_{0}$ and $B_{1}$ only along with the fact that $\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle$.
@@
$\hat{E}_{1}=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/\lambda_{\theta,\phi}^{2}$.
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)\tilde{X}_{B}-\cos(2\theta)\tilde{Z}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{5}$, where $\delta_{5}:=\delta_{0}/\lambda_{\theta,\phi}^{2}+\sin^{2}(2\theta)(\delta_{4}+\delta_{\mathsf{QHE}})+\cos^{2}(2\theta)(\delta_{0}+\delta_{\mathsf{QHE}})$.
@@
$\leq\big{\|}\big{(}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\sin(2\theta)\big{(}X_{B}-\tilde{X}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}\cos(2\theta)\big{(}Z_{B}-\tilde{Z}_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Next we take the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$, and sum over $\alpha$. This allows us to apply 8 to the first term, Lemma 3.1 followed by 6 to the second, and Lemma 3.1 followed by 5 to the final term, giving the desired bound. ∎
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{6}$, where $\delta_{6}:=\delta_{0}+\delta_{\mathsf{QHE}}+2(\delta_{4}+\delta_{\mathsf{QHE}})+(\delta_{0}+\delta_{\mathsf{QHE}})/(2\sin^{2}(\phi))+(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2}+(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\delta_{\mathsf{QHE}})$.
@@
$\big{\|}\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\big{\|}\tilde{X}_{B}\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
@@
For the first term, we can use the fact that $\tilde{X}_{B}$ is unitary, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1 to obtain a bound of $\delta_{0}+\delta_{\mathsf{QHE}}$. For the second term,
$\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\big{\|}(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
@@
Since $\tilde{Z}_{B}$ is unitary $\big{\|}\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\leq 2$, hence when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 6 to $\|(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\|^{2}$; the product results in a contribution $2(\delta_{4}+\delta_{\mathsf{QHE}})$. Next, we note that
$\big{\|}\big{(}Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}X_{B}\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|X_{B}\|_{\mathrm{op}}^{2}\|\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
@@
where the equality follows since $\{Z_{B},X_{B}\}=0$ by definition. Now $\|X_{B}\|_{\mathrm{op}}^{2}\leq 1/(2\sin^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1, resulting in a contribution of $(\delta_{0}+\delta_{\mathsf{QHE}})/(2\sin^{2}(\phi))$.
We now address the term $\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$. 8 tells us that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big{\|}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2},$ (4.16)
where $T_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)Z_{B})$, which implies
$\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}X_{B}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}\tilde{Z}_{B}-Z_{B}\big{\|}_{\mathrm{op}}^{2}\big{\|}\big{(}X_{B}-(-1)^{\mathsf{Dec}(\alpha)}T_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
@@
We have that $\big{\|}\tilde{Z}_{B}-Z_{B}\big{\|}_{\mathrm{op}}^{2}\leq 1+1/(2\cos^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Eq. 4.16 to obtain a contribution of $(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\lambda_{\theta,\phi})^{2}$. Finally, we bound the last term using the fact that $\tilde{Z}_{B}$ and $Z_{B}$ commute, that is
$\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}T_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\big{\|}T_{\theta,\phi}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\|T_{\theta,\phi}\|^{2}_{\mathrm{op}}\big{\|}\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
@@
The operator norm is bounded by $1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Lemma 3.1 followed by 5, resulting in an overall contribution of $(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\delta_{\mathsf{QHE}})$. By summing all the contributions from every use of the triangle inequality, we obtain the desired bound. ∎
@@
where $\delta_{7}:=\delta_{6}/4+\delta_{5}/(2\sin^{2}(2\theta))$.
@@
Consider the choice $|\mathsf{aux}_{\alpha|\chi}\rangle=\frac{1}{\cos(\theta)}P_{B}^{0}|\Psi_{\alpha|\chi}\rangle$. Then
[…]
@@
$\leq\big{\|}\big{(}P_{B}^{0}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\big{\|}(\tilde{X}_{B}P_{B}^{1}-P_{B}^{0}\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
Taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term by $\delta_{6}/4$ using 11. Using the triangle inequality again,
[…]
$\leq\big{\|}\big{(}(-1)^{\mathsf{Dec}(\alpha)}P_{B}^{0}\tilde{T}_{\theta,\phi}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}P_{B}^{0}\big{(}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq\big{\|}\big{(}P_{B}^{0}\tilde{T}_{\theta,\phi}-\tan(\theta)P_{B}^{0}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}P_{B}^{0}\big{\|}_{\mathrm{op}}^{2}\big{\|}\big{(}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
where $\tilde{T}_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)\tilde{Z}_{B})$. Recall that $\big{\|}P_{B}^{0}\big{\|}_{\mathrm{op}}^{2}\leq 1/2$, and after taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term using 9, resulting in a contribution of $\delta_{5}/(2\sin^{2}(2\theta))$ to the final bound. Finally, by direct calculation we obtain
@@
Eqs. 4.15 and 4.17 complete the self-testing statement of the model without the action of measurements in the second round according to Eq. 3.21.
Extracting the measurements: Next we turn our attention to the measurements.
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}(VZ_{B}-(\sigma_{Z}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{8,x}$, where $\delta_{8,x}=\delta_{0}+x\cdot\delta_{\mathsf{QHE}}$ for $x\in\{0,1\}$.
@@
For the second, we used the fact that $\|P_{B}^{b}\|_{\mathrm{op}}^{2}\leq 1/2$ and $\|\tilde{X}_{B}P_{B}^{1}\|_{\mathrm{op}}=\|P_{B}^{1}\|_{\mathrm{op}}$ since $\tilde{X}_{B}$ is unitary. For the case $\mathsf{Enc}(x=0)=\chi$, taking the expectation, summing over $\alpha$ and applying 5 recovers the desired bound. For the case $\mathsf{Enc}(x=1)=\chi$, the same procedure with an intermediate application of Lemma 3.1, which accumulates an extra factor of $\delta_{\mathsf{QHE}}$, obtains the desired bound. ∎
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{9,x}$, where $\delta_{9,x}=\delta_{4}+\delta_{\mathsf{QHE}}\cdot(2-x)+\delta_{6}/2$.

### v3

@@ 4.2 Compiled self-test for tilted-CHSH inequalities
We now proceed to study the relations imposed by the second SOS polynomial, which imply $\hat{E}_{1}\leq\delta_{0}/\tau_{\theta,\phi}^{2}$. Before we begin, we note that in any earlier claim that did not include a factor of $(-1)^{\mathsf{Dec}(\alpha)}$, we can always apply Lemma 3.1 to exchange the expectation over encryptions of $x=0$ to an expectation over encryptions of $x=1$ at the cost of adding $\mathrm{negl}(\lambda)$ to the bound. This follows from the fact that such claims are bounding the expectation of operators $\mathcal{N}$ constructed from sums of products of $B_{0}$ and $B_{1}$ only along with the fact that $\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle\approx_{\mathrm{negl}(\lambda)}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\mathcal{N}|\Psi_{\alpha|\chi}\rangle$.
@@
$\hat{E}_{1}=\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big\|\big(\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{0}/\tau_{\theta,\phi}^{2}$.
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big\|\big(\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)\tilde{X}_{B}-\cos(2\theta)\tilde{Z}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{5}$, where $\delta_{5}:=2\delta_{0}/\tau_{\theta,\phi}^{2}+4\sin^{2}(2\theta)(\delta_{4}+\mathrm{negl}(\lambda))+4\cos^{2}(2\theta)(\delta_{0}+\mathrm{negl}(\lambda))$.
@@
$\big\|\big(\mathbbm{1}$- $(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)\tilde{X}_{B}-\cos(2\theta)\tilde{Z}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|$
$\leq\big\|\big(\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|$
$+\big\|\sin(2\theta)\big(X_{B}-\tilde{X}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|+\big\|\cos(2\theta)\big(Z_{B}-\tilde{Z}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|.$
Next, we square both sides, and use the fact that, for three real numbers $a,b,c$, $(a+b+c)^{2}\leq 4(a^{2}+b^{2})+2c^{2}$, to arrive at
@@
$\leq 2\big\|\big(\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+4\big\|\sin(2\theta)\big(X_{B}-\tilde{X}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}+4\big\|\cos(2\theta)\big(Z_{B}-\tilde{Z}_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$
We now take the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$, and sum over $\alpha$. This allows us to apply 8 to the first term, Lemma 3.1 followed by 6 to the second, and Lemma 3.1 followed by 5 to the final term, giving the desired bound. ∎
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big\|\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{6}$, where $\delta_{6}:=2(\delta_{0}+\mathrm{negl}(\lambda))+8(\delta_{4}+\mathrm{negl}(\lambda))+8(\delta_{0}+\mathrm{negl}(\lambda))/(2\sin^{2}(\phi))+16(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\tau_{\theta,\phi})^{2}+16(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\mathrm{negl}(\lambda))$.
@@
$\big\|\{\tilde{Z}_{B},\tilde{X}_{B}\}|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq 2\big\|\tilde{X}_{B}\big(\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}+2\big\|\big(\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$
@@
For the first term, we can use the fact that $\tilde{X}_{B}$ is unitary, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1 to obtain a bound of $2(\delta_{0}+\mathrm{negl}(\lambda))$. For the second term,
$2\big\|\big(\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)\tilde{X}_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}$ $\leq 4\big\|\big(\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+4\|\big(\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)X_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$\leq 4\big\|\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\big\|(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+8\|\big(Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)X_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+8\|\big(\tilde{Z}_{B}-Z_{B}\big)X_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$
@@
Since $\tilde{Z}_{B}$ is unitary $\big\|\tilde{Z}_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\|_{\mathrm{op}}^{2}\leq 2$, hence when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$ for the first term above, we can apply Lemma 3.1 followed by 6 to $\|(\tilde{X}_{B}-X_{B})|\Psi_{\alpha|\chi}\rangle\|^{2}$; the product results in a contribution $8(\delta_{4}+\mathrm{negl}(\lambda))$. Next, we note that
$8\big\|\big(Z_{B}+(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)X_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}$ $=8\big\|X_{B}\big(Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$\leq 8\|X_{B}\|_{\mathrm{op}}^{2}\|\big(Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2},$
@@
where the equality follows since $\{Z_{B},X_{B}\}=0$ by definition. Now $\|X_{B}\|_{\mathrm{op}}^{2}\leq 1/(2\sin^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and sum over $\alpha$, we can apply Lemma 3.1 followed by 1, resulting in a contribution of $8(\delta_{0}+\mathrm{negl}(\lambda))/(2\sin^{2}(\phi))$.
We now address the term $8\|\big(\tilde{Z}_{B}-Z_{B}\big)X_{B}|\Psi_{\alpha|\chi}\rangle\big\|^{2}$. 8 tells us that
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\big\|\big(X_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tau_{\theta,\phi}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{0}/(\sin(2\theta)\tau_{\theta,\phi})^{2},$ (4.11)
where $\tau_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)Z_{B})$, which implies
[…]
@@
For the second term above, we have that $\big\|\tilde{Z}_{B}-Z_{B}\big\|_{\mathrm{op}}^{2}\leq 1+1/(2\cos^{2}(\phi))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Eq. 4.11 to obtain a contribution of $16(1+1/(2\cos^{2}(\phi)))\delta_{0}/(\sin(2\theta)\tau_{\theta,\phi})^{2}$. Finally, we bound the first term above using the fact that $\tilde{Z}_{B}$ and $Z_{B}$ commute, that is,
$6\|\big(\tilde{Z}_{B}-Z_{B}\big)\tau_{\theta,\phi}|\Psi_{\alpha|\chi}\rangle\big\|^{2} =1$$=6\big\|\tau_{\theta,\phi}\big(\tilde{Z}_{B}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
[…]
@@
The operator norm is bounded by $1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta))$, and when taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can apply Lemma 3.1 followed by 5, resulting in a contribution of $16(1/\sin^{2}(2\theta)+1/(2\cos^{2}(\phi)\tan^{2}(2\theta)))(\delta_{0}+\mathrm{negl}(\lambda))$. By summing all the contributions from every use of the triangle inequality, we obtain the desired bound. ∎
@@
where $\delta_{7}:=\delta_{6}/2+2\delta_{5}/(\sin^{2}(2\theta))$.
@@
Consider the choice $|\mathsf{aux}_{\alpha|\chi}\rangle=\frac{1}{\cos(\theta)}P_{B}^{0}|\Psi_{\alpha|\chi}\rangle$. Then,
$\big\|$ $V|\Psi_{\alpha|\chi}\rangle-\big(\cos(\theta)|0\rangle+(-1)^{\mathsf{Dec}(\alpha)}\sin(\theta)|1\rangle\big)\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2}$
@@
$\leq 2\big\|\big(P_{B}^{0}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}+2\big\|(\tilde{X}_{B}P_{B}^{1}-P_{B}^{0}\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$
Taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term by $\delta_{6}/2$ using 11. Using the triangle inequality again,
$2\big\|$ $\big(P_{B}^{0}\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$\leq 4\big\|\big((-1)^{\mathsf{Dec}(\alpha)}P_{B}^{0}\tilde{T}_{\theta,\phi}-(-1)^{\mathsf{Dec}(\alpha)}\tan(\theta)P_{B}^{0}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+4\big\|P_{B}^{0}\big(\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$\leq 4\big\|\big(P_{B}^{0}\tilde{T}_{\theta,\phi}-\tan(\theta)P_{B}^{0}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+4\big\|P_{B}^{0}\big\|_{\mathrm{op}}^{2}\big\|\big(\tilde{X}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\tilde{T}_{\theta,\phi}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2},$
where $\tilde{T}_{\theta,\phi}=\frac{1}{\sin(2\theta)}(\mathbbm{1}-\cos(2\theta)\tilde{Z}_{B})$. Recall that $\big\|P_{B}^{0}\big\|_{\mathrm{op}}^{2}\leq 1/2$, and after taking the expectation over $\chi:\mathsf{Enc}(x=1)=\chi$ and summing over $\alpha$, we can bound the second term using 9, resulting in a contribution of $2\delta_{5}/(\sin^{2}(2\theta))$ to the final bound. Finally, by direct calculation we obtain
@@
Eqs. 4.10 and 4.12 complete the self-testing statement of the model without the action of measurements in the second round according to Definition 4.5.
Extracting the measurements
Next we turn our attention to the measurements.
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big\|(VZ_{B}-(\sigma_{Z}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{8,x}$, where $\delta_{8,x}=\delta_{0}+x\cdot\mathrm{negl}(\lambda)$ for $x\in\{0,1\}$.
@@
For the second inequality, we used the fact that $\|P_{B}^{b}\|_{\mathrm{op}}^{2}\leq 1/2$ and $\|\tilde{X}_{B}P_{B}^{1}\|_{\mathrm{op}}=\|P_{B}^{1}\|_{\mathrm{op}}$ since $\tilde{X}_{B}$ is unitary. For the case $\mathsf{Enc}(x=0)=\chi$, taking the expectation, summing over $\alpha$ and applying 5 recovers the desired bound. For the case $\mathsf{Enc}(x=1)=\chi$, the same procedure with an intermediate application of Lemma 3.1, which accumulates an extra factor of $\mathrm{negl}(\lambda)$, obtains the desired bound. ∎
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big\|(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{9,x}$, where $\delta_{9,x}=2\delta_{4}+2\mathrm{negl}(\lambda)\cdot(2-x)+\delta_{6}$.